The area of a rectangle is 49in. Express the perimeter as a function of the width w, and state the domain of this function.
I answer it this way. 98= 2(L+W)
98/2 = 2(L+W)/2
49 = (L+W)
49-W = L+W-W
0 < W <49

Answer by chessace(471)

Wow, how did you start like that?
A=L*W=49
P=2(L+W)=2(49/W + W)
P=98/W + 2W answers the first part.
The only problem is /0 so W>0 answers the 2nd part.